OpenStudy (anonymous):

Use L'Hopital's Rule to calculate: lim e^3t -1 ------- (e-1, e is raised to 3t) t t ->0

OpenStudy (amistre64):

whats that stuff in (..)

OpenStudy (anonymous):

just explaining what the numerator is, in case anyone was confused.

OpenStudy (amistre64):

ok, and the other question i have is; if there is no x to limit in this, where do plug in the values for x at?>

OpenStudy (anonymous):

opps its suppose to be as x goes to 0, not t. sorry

OpenStudy (amistre64):

ok, x to 0 the value of t as x goes to 0 is t the value of e^-3t -1 as x goes to 0 is e^-3t-1

OpenStudy (amistre64):

same concept but with the e^3t lol

OpenStudy (amistre64):

Lhop is used for indeterminant conditions, such as 0/0 and inf/inf

OpenStudy (amistre64):

lol, i see you change x to t now

OpenStudy (anonymous):

yea lol it was suppose to be t not x.

OpenStudy (amistre64):

ok, so lets take the derivative of the top and the derivative of the bottm; and ratio them so see how fast they are moving during any given point

OpenStudy (amistre64):

derivate of t wrt t is?

OpenStudy (anonymous):

what is "wrt"?

OpenStudy (amistre64):

its the lazy form of "With Respect To"

OpenStudy (anonymous):

ooo ok lol. so t would be 0 or 1. i think it's 0 tho.

OpenStudy (amistre64):

\[\frac{d}{dt}t\to\frac{dt}{dt}=1\]

OpenStudy (amistre64):

now for the harder one :) what is e^3t -1 wrt t?

OpenStudy (anonymous):

lol. 3e^2t??

OpenStudy (amistre64):

very close; except for bases the exponent stays put. \[D[x^n]=nx^{n-1}\] \[D[B^{nx}]=nB^{nx}+ln(B)\]

OpenStudy (amistre64):

not + but *

OpenStudy (amistre64):

so\[D[e^3{t}]\to3e^{3t}ln(e)=3e^{3t}\]

OpenStudy (amistre64):

if i could work the latex code correctly that would be more impressive; as is you kinda gotta use your imagination :)

OpenStudy (anonymous):

lol. that was easier than i thought! thanks for doing it step by step with me. :)

OpenStudy (amistre64):

so the limit of 3e^(3t) as t moves to 0 is: 3

OpenStudy (amistre64):

yw